Find the most general antiderivative of the function. Use C for any needed constant. (Check your answer by differentiation.) (x>0)
g(x) = (6 - 4 x^3 + 3 x^(7))/x^(7)

G(x)=
1
Expert's answer
2011-03-17T07:01:31-0400
g(x) = 6/x7 - 4x3/x7 + 3x7/x7 = 6x-7 - 4x-4 + 3

G(x) = int(g(x), dx) = -6/6 x-6 + 4/3 x-3 + 3x + C = - x-6 +4/3 x-3 + 3x + C